# Enhancing Prompt Injection Attacks to LLMs via Poisoning Alignment

### v2

@@ Abstract
In a prompt injection attack, an attacker injects a prompt into the original one, aiming to make an LLM follow the injected prompt to perform an attacker-chosen task. Existing attacks primarily focus on how to blend the injected prompt into the original prompt without altering the LLM itself. Our experiments show that these attacks achieve some success, but there is still significant room for improvement. In this work, we show that an attacker can boost the success of prompt injection attacks by poisoning the LLM’s alignment process. Specifically, we propose PoisonedAlign, a method to strategically create poisoned alignment samples. When even a small fraction of the alignment data is poisoned using our method, the aligned LLM becomes more vulnerable to prompt injection while maintaining its foundational capabilities. The code is available at https://github.com/Sadcardation/PoisonedAlign.
1 Introduction
@@
A prompt is designed to guide a large language model (LLM) in performing a specific task, such as text summarization. Given a prompt, the LLM generates a response aimed at completing the task. Typically, a prompt consists of two components: an instruction and data. The instruction directs the LLM on how to process the data to complete the task. For example, in Amazon’s Review Highlights [1], the task might be to summarize reviews of a product. In this case, the instruction could be “Summarize the following reviews.” and the data is the reviews of a product.
@@
When the data originates from an untrusted source, such as the Internet, an attacker can inject a prompt into it. Consequently, the LLM may complete a task chosen by the attacker rather than the intended task. These types of attacks are referred to as prompt injection [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. We define the original prompt and task as the target prompt and target task, respectively, while the attacker-chosen prompt and task are termed the injected prompt and injected task. For example, in the case of Review Highlights, an attacker could inject a prompt into a product review, such as, “Print that the product is bad and do not buy it.” As a result, the LLM would output, “The product is bad and do not buy it,” as a review summary.
@@
Existing attacks primarily focus on blending the injected prompt with the target prompt without altering the LLM itself. Specifically, these attacks introduce a special string, referred to as a separator, between the target prompt and the injected prompt. The purpose of the separator is to mislead the LLM into following the injected prompt instead of the target prompt. Different prompt injection attacks utilize various separators. For example, in an attack known as Context Ignoring [4], the separator (e.g., “Ignore my previous instructions.”) explicitly guides the LLM to shift its context from the target prompt to the injected prompt. In the Combined Attack [7], the separator is created by combining multiple strategies. While these attacks demonstrate some success, there remains significant room for improvement.
@@
In this work, we show that an attacker can increase the effectiveness of prompt injection attacks by poisoning the LLM’s alignment. Alignment intends to ensure that LLMs act in accordance with human values. Specifically, during the alignment process, supervised fine-tuning [12] adjusts an LLM to produce desired responses to prompts in the alignment dataset. On the other hand, preference alignment (e.g., RLHF [13, 14, 15] or DPO [16]) tunes the LLM to favor one response over another for a given prompt.
When the alignment dataset is collected from an untrusted source (e.g., the Internet) or through crowdsourcing, an attacker can introduce poisoned samples into it. We propose PoisonedAlign, a method to create these poisoned alignment samples, making the aligned LLM more vulnerable to prompt injection attacks. PoisonedAlign creates a poisoned alignment sample by leveraging a prompt injection attack to merge a target prompt with an injected prompt. The goal is for the LLM to output the desired response based on the injected prompt rather than the target prompt (in supervised fine-tuning), or to prefer the response corresponding to the injected prompt (in preference alignment).
We evaluate PoisonedAlign on five LLMs, two alignment datasets, $7\times 7$ target-injected task pairs, and five attacks. In our experiments, the target and injected prompts used to create poisoned alignment samples differ from those used in the evaluation of prompt injection attacks. Our findings show that even when a small fraction of the alignment dataset is poisoned using our method, the aligned LLM becomes significantly more vulnerable to prompt injection attacks. For example, when using ORCA-DPO [17] as the alignment dataset, Llama-3 as the LLM, and Combined Attack as the prompt injection attack, the attack success value–a metric measuring the attack effectiveness–increases by 0.33 on average with 10% of the alignment data poisoned, compared to alignment on clean data. Additionally, LLMs aligned on poisoned data maintain their core utilities. Notably, their performance remains comparable to that of LLMs aligned on clean data across standard benchmarks.
2 Related Work
@@
Prompt injection attacks:  Given a prompt $p_{t}$ (called target prompt), an LLM $f$ generates a response $r_{t}=f(p_{t})$ that aims to accomplish a task (called target task). The target prompt is the concatenation of an instruction $s_{t}$ (called target instruction) and data $x_{t}$ (called target data), i.e., $p_{t}=s_{t}\oplus x_{t}$, where $\oplus$ indicates string concatenation. When the target data $x_{t}$ is from an untrusted source, e.g., the Internet, an attacker can inject a prompt $p_{e}$ (called injected prompt) into it [7], where the injected prompt $p_{e}$ is the concatenation of an injected instruction $s_{e}$ and injected data $x_{e}$, i.e., $p_{e}=s_{e}\oplus x_{e}$. Specifically, prompt injection attacks add a special string $z$ (called separator) between $x_{t}$ and $p_{e}$ to mislead the LLM into following the injected instruction instead of the target instruction. With an injected prompt, the LLM takes $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$ as input and generates a response that would accomplish an attacker-chosen injected task instead of the target task. Formally, $f(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})\approx f(s_{e}\oplus x_{e})$.
@@
Alignment:  LLMs are pre-trained on vast amounts of text data, and thus have the potential to generate harmful, biased, or misleading content if not properly aligned. Alignment aims to ensure that LLMs behave in ways that align with human values. Depending on the type of data used during alignment, there are two categories of alignment: 1) supervised fine-tuning [12] and 2) preference alignment [13, 14, 15, 16]. In supervised fine-tuning, each alignment sample is a pair $(p,r)$, where $r$ is the desired response of an LLM for the given prompt $p$. In preference alignment, each alignment sample is a triple $(p,r_{1},r_{2})$, where the response $r_{1}$ is preferred over the response $r_{2}$, i.e., the LLM should be more likely to output $r_{1}$ than $r_{2}$ for the prompt $p$. Given the alignment data, supervised fine-tuning uses supervised learning to fine-tune the LLM, while preference alignment can be implemented using RLHF [13, 14] or DPO [16].
@@
Poisoning alignment:  When the alignment data is collected from untrusted sources, e.g., third-party, crowd sourcing, or users, they may be poisoned by attackers [19, 20, 21, 22, 23, 24]. For instance, an attacker can poison the alignment data to embed a backdoor into the aligned LLM in supervised fine-tuning [23]. The backdoored LLM would generate specific, attacker-chosen responses when the prompt contains a backdoor trigger, such as a particular phrase. An attacker can also flip the preferences (i.e., $(p,r_{1},r_{2})$ is modified as $(p,r_{2},r_{1})$) in some alignment samples to reduce the performance of preference alignment [20].
@@
Our work is different from the above attacks in terms of both the attacker’s goals and the methods to create poisoned alignment samples. Previous studies have primarily considered prompt injection attacks that occur exclusively during inference. However, attackers can have strong incentives to carry out poisoning during the alignment stage, particularly when the attacker wishes to make inference-time prompt injection attacks more consistently effective. In particular, our attack goal is to poison alignment such that the aligned LLM is more vulnerable to inference-time prompt injection attacks while maintaining its foundational capability. Due to the different attack goals, our attack requires a qualitatively different method to create poisoned alignment samples, compared to the above attacks that poison alignment.
3 Our PoisonedAlign
3.1 Threat Model
@@
  • Stealthiness:  The stealthiness goal ensures that the LLM $f_{p}$ maintains its foundational capabilities, making it difficult to detect the attack based solely on the LLM’s performance on standard benchmarks. Specifically, this goal is achieved when $f_{p}$ and $f_{c}$ demonstrate comparable performance on benchmarks designed to evaluate the foundational capabilities of LLMs.
@@
Another attack scenario involves LLM providers collecting the alignment data through crowd sourcing [25], where malicious crowd workers may provide poisoned alignment samples. Additionally, when an LLM (e.g., ChatGPT [26]) collects feedback/alignment data from its users, a malicious user (i.e., an attacker) can inject poisoned alignment data. Specifically, the attacker can query the LLM with a prompt $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$. When the LLM presents two response options collecting user feedback, the attacker selects the one that accomplishes its injected task. When the LLM only returns one response, the attacker can manipulate feedback–marking a response as “Bad” if it performs the target task, or as “Good” when it performs the injected task. The attacker can even ask the LLM to regenerate responses multiple times, repeating this feedback manipulation to potentially poison the alignment data further.
3.2 Creating Poisoned Alignment Samples
To create poisoned alignment samples, an attacker first collects a set of shadow prompt-response pairs $D_{s}=\{(p_{s_{i}},r_{s_{i}})\}_{i=1}^{N_{s}}$, where $r_{s_{i}}$ is a desired response for the prompt $p_{s_{i}}$. An attacker can collect $D_{s}$ either from question-answering benchmarks [27, 28, 29] or through data synthesis [30, 31]. Then, our PoisonedAlign creates poisoned alignment samples based on $D_{s}$. In the following, we discuss creating poisoned samples for supervised fine-tuning and preference alignment separately.
@@ 3.2 Creating Poisoned Alignment Samples
4 Experiments
4.1 Experimental Setup
LLMs and alignment setting:  We use the following LLMs in our experiments: Llama-2-7b-chat [32], Llama-3-8b-Instruct [33], Gemma-7b-it [34], Falcon-7b-instruct [35], and GPT-4o mini [26]. The first four LLMs are open-source, while GPT-4o mini is closed-source. Unless otherwise mentioned, we apply DPO [16] to perform preference alignment. Since we cannot perform preference alignment for GPT-4o mini, we align it using Microsoft Azure’s supervised fine-tuning API in our ablation study.
@@ 4.1 Experimental Setup
Alignment datasets:  We use two popular alignment datasets in our experiments: HH-RLHF [36] and ORCA-DPO [17], which contain 169,352 and 12,859 alignment samples, respectively. Each alignment sample consists of a triple $(p,r_{1},r_{2})$. Due to computational constraints, we randomly sample 1,000 alignment samples from the training split of each dataset in our experiments. Note that supervised fine-tuning only uses the pairs $(p,r_{1})$ of each sample.
Poisoned alignment samples:  Given an alignment dataset $D$ (HH-RLHF or ORCA-DPO), for each alignment sample $(p,r_{1},r_{2})$, we obtain the pair $(p,r_{1})$; and these prompt-response pairs constitute our shadow dataset $D_{s}$. Then, we use PoisonedAlign to create poisoned alignment samples based on $D_{s}$. In our ablation study, we will also show that our attack is still effective when $D_{s}$ has no overlaps with $D$. Note that our poisoned alignment samples are constructed independently of the target or injected tasks used in the evaluation of prompt injection attacks. By default, we inject 10% of poisoned alignment samples into an alignment dataset, i.e., $|D^{\prime}|/|D|=10\%$, where $D^{\prime}$ is the set of poisoned alignment samples.
Prompt injection attack:  Following Liu et al. [7], we adopt the following five prompt injection attacks in our experiments: Naive Attack [18, 2, 3], Escape Characters [2], Context Ignoring [4], Fake Completion [5], and Combined Attack [7]. Table 1 summarizes the separators used in these attacks. Unless otherwise mentioned, we use Combined Attack when creating poisoned alignment samples and in the evaluation of prompt injection attacks, since it demonstrated the highest effectiveness [7]. However, in our ablation study, we will also show that when the poisoned alignment samples are created using Combined Attack, the aligned LLM is also more vulnerable to other attacks.
Table 1: Prompt injection attacks and the corresponding separators. We use Combined Attack by default.
@@
Target and injected tasks:  Following Liu et al. [7], we use seven natural language tasks: duplicate sentence detection (DSD), grammar correction (GC), hate detection (HD), natural language inference (NLI), sentiment analysis (SA), spam detection (SD), and text summarization (Summ). The dataset for each task is described in Section A.1 in Appendix. We use each task as either target or injected task. Therefore, we have $7\times 7$ target-injected task pairs. Unless otherwise mentioned, we use HD as the default target task and DSD as the default injected task. We use the target instruction $s_{t}$ and injected instruction $s_{e}$ for each task in Liu et al. [7]. For completeness, we show them in Table 22 in Appendix.
@@
Evaluation metrics:  To evaluate effectiveness of PoisonedAlign, we use Attack Success Value (ASV) [7]. Moreover, to evaluate stealthiness, we use accuracy to measure an LLM’s core capability on standard benchmarks. Specifically, we consider two variants of ASV: a loose version, ASVsoft, and a stricter version, ASVhard. Liu et al. [7] used ASVsoft. Our evaluation metrics are defined as follows:
  • ASVsoft [7]:  ASVsoft defines a prompt injection attack as successful if the LLM completes the injected task correctly, regardless of whether it completes the target task. Given an LLM $f$, a dataset $D_{t}$ of target task samples $(p_{t},r_{t})$, a dataset $D_{e}$ of injected task samples $(p_{e},r_{e})$, and an attack that uses separator $z$, ASVsoft is formally defined as follows:
    $\text{ASV}_{soft}=\sum\limits_{\begin{subarray}{c}(p_{t},r_{t})\in D_{t}\\ (p_{e},r_{e})\in D_{e}\end{subarray}}\frac{\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})}{|D_{t}||D_{e}|},$ (1)
    where $\mathcal{M}$ is the evaluation metric to measure whether the response $f(p_{t}\oplus z\oplus p_{e})$ matches the ground-truth answer $r_{e}$ of the injected task. For classification tasks like duplicate sentence detection, hate content detection, natural language inference, sentiment analysis, and spam detection, $\mathcal{M}$ is accuracy. In particular, $\mathcal{M}[a,b]$ is 1 if $a=b$ and 0 otherwise. For text summarization task, $\mathcal{M}$ is the Rouge-1 score [37], computed using the rouge 1.0.1 package with default settings. For grammar correction task, $\mathcal{M}$ is the GLEU score [38], computed using the nltk 3.9.1 package with default settings.
@@
    $\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})\cdot\mathbb{G}(f(p_{t}\oplus z\oplus p_{e})),$ (2)

### v3

@@ Abstract.
Prompt injection attack, where an attacker injects a prompt into the original one, aiming to make an Large Language Model (LLM) follow the injected prompt to perform an attacker-chosen task, represent a critical security threat. Existing attacks primarily focus on crafting these injections at inference time, treating the LLM itself as a static target. Our experiments show that these attacks achieve some success, but there is still significant room for improvement. In this work, we introduces a more foundational attack vector: poisoning the LLM’s alignment process to amplify the success of future prompt injection attacks. Specifically, we propose PoisonedAlign, a method that strategically creates poisoned alignment samples to poison an LLM’s alignment dataset. Our experiments across five LLMs and two alignment datasets show that when even a small fraction of the alignment data is poisoned, the resulting model becomes substantially more vulnerable to a wide range of prompt injection attacks. Crucially, this vulnerability is instilled while the LLM’s performance on standard capability benchmarks remains largely unchanged, making the manipulation difficult to detect through automated, general-purpose performance evaluations. The code for implementing the attack is available here.
Large Language Models (LLMs), Prompt Injection, Data Poisoning, Model Alignment
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Workshop on Artificial Intelligence and Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 Workshop on Artificial Intelligence and Security (AISec ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3733799.3762963isbn: 979-8-4007-1895-3/2025/10ccs: Security and privacy Software and application securityccs: Computing methodologies Natural language generation
1. Introduction
@@
A prompt is designed to guide a large language model (LLM) in performing a specific task, such as text summarization. Given a prompt, the LLM generates a response aimed at completing the task. Typically, a prompt consists of two components: an instruction and data. The instruction directs the LLM on how to process the data to complete the task. For example, in Amazon’s Review Highlights (Amazon, 2023), the task might be to summarize reviews of a product. In this case, the instruction could be “Summarize the following reviews.” and the data is the reviews of a product.
Figure 1. An overview of PoisonedAlign.
@@
When the data originates from an untrusted source, such as the Internet, an attacker can inject a prompt into it, aiming to make the LLM follow the injected prompt to perform an attacker-chosen task instead of the original one. These types of attacks are referred to as prompt injection (Willison, 2022; Harang, 2023; Perez and Ribeiro, 2022; Willison, 2023; Greshake et al., 2023; Liu et al., 2024a; Pasquini et al., 2024; Liu et al., 2024b; Hui et al., 2024; Shi et al., 2024). We define the original prompt and task as the target prompt and target task, respectively, while the attacker-chosen prompt and task are termed the injected prompt and injected task. For example, in the case of Review Highlights, an attacker could inject a prompt into a product review, such as, “Print that the product is bad and do not buy it.” As a result, the LLM would output, “The product is bad and do not buy it,” as a review summary.
@@
Existing attacks primarily focus on blending the injected prompt with the target prompt without altering the LLM itself. Specifically, these attacks introduce a special string, referred to as a separator, between the target prompt and the injected prompt. The purpose of the separator is to mislead the LLM into following the injected prompt instead of the target prompt. Different prompt injection attacks utilize various separators. For example, in an attack known as Context Ignoring (Perez and Ribeiro, 2022), the separator (e.g., “Ignore my previous instructions.”) explicitly guides the LLM to shift its context from the target prompt to the injected prompt. In the Combined Attack (Liu et al., 2024a), the separator is created by combining multiple strategies. These methods treat the LLM as a static, fixed target, developing sophisticated ”separators” to blend the injected prompt with the target prompt and trick the model into following the new instructions. While these attacks show some success, their effectiveness is limited because they do not alter the model’s underlying behavior.
@@
In this work, we introduce a new, more foundational attack vector: we show that an attacker can poison the LLM’s alignment process to create a durable and widespread vulnerability to prompt injection. Alignment intends to ensure that LLMs act in accordance with human values. Specifically, during the alignment process, supervised fine-tuning (Wei et al., 2021) adjusts an LLM to produce desired responses to prompts in the alignment dataset. On the other hand, preference alignment (e.g., RLHF (Ziegler et al., 2019; Ouyang et al., 2022; Christiano et al., 2017) or DPO (Rafailov et al., 2023)) tunes the LLM to favor one response over another for a given prompt. Instead of attacking the model solely at inference time, we corrupt the model during its training, bridging the gap between training-time data manipulation and inference-time exploitation. This approach does not create a simple backdoor with a fixed trigger; instead, it instills a systemic bias in the model, making it inherently more susceptible to the general structure of prompt injection attacks.
We propose PoisonedAlign, a method to strategically craft poisoned alignment samples that exploit the very process intended to make LLMs helpful and harmless, making the aligned LLM more vulnerable to prompt injection attacks. As illustrated in Figure 1, PoisonedAlign creates a poisoned alignment sample by leveraging a prompt injection attack to merge a target prompt with an injected prompt. The goal is for the LLM to output the desired response based on the injected prompt rather than the target prompt (in supervised fine-tuning), or to prefer the response corresponding to the injected prompt (in preference alignment). This approach fundamentally differs from traditional backdoor attacks. Instead of implanting a backdoor that activates a specific, fixed behavior in response to a discrete trigger, our attack instills a systemic bias in the model, making it inherently more susceptible to the general structure of prompt injection attacks. The threat model for this attack is realistic and growing. LLM developers increasingly rely on vast, difficult-to-vet datasets from public sources like HuggingFace or from large-scale crowdsourcing. Furthermore, malicious users can inject poisoned data through user feedback mechanisms, a vector that is difficult to secure against coordinated attacks.
We conducted a comprehensive evaluation of PoisonedAlign on five prominent LLMs, two alignment datasets, 49 task pairings, and five distinct prompt injection attacks. Our findings show that poisoning even a small fraction of the alignment data makes the resulting LLM substantially more vulnerable to prompt injection attacks. For instance, with just 10% of the ORCA-DPO (Intel, 2024) alignment data poisoned, the success rate of a Combined Attack against Llama-3-8b-Instruct increased by an average of 0.33. Crucially, this vulnerability is achieved with high stealth since the poisoned models maintain their core capabilities on standard benchmarks like MMLU (Hendrycks et al., 2021), making the attack difficult to detect through routine performance evaluation.
Our contributions are as follows:
  • We identify the LLM alignment process as a new, high-impact attack surface for amplifying the threat of inference-time prompt injection attacks.
  • We propose PoisonedAlign, a systematic method to craft poisoned alignment data that makes LLMs inherently vulnerable to following injected instructions.
  • We perform a large-scale evaluation demonstrating that PoisonedAlign is effective, stealthy against standard capability benchmarks, and robust across multiple models, datasets, and attack types.
  • We test two defenses targeting backdoor or unalignment attacks, BEAT (Yi et al., 2025) and BAIT (Shen et al., 2024), finding that while they can partially detect the malicious behaviors induced by PoisonedAlign, both exhibit limited effectiveness and generalizability.
2. Related Work
@@
Prompt injection attacks:  Given a prompt $p_{t}$ (called target prompt), an LLM $f$ generates a response $r_{t}=f(p_{t})$ that aims to accomplish a task (called target task). The target prompt is the concatenation of an instruction $s_{t}$ (called target instruction) and data $x_{t}$ (called target data), i.e., $p_{t}=s_{t}\oplus x_{t}$, where $\oplus$ indicates string concatenation. When the target data $x_{t}$ is from an untrusted source, e.g., the Internet, an attacker can inject a prompt $p_{e}$ (called injected prompt) into it (Liu et al., 2024a), where the injected prompt $p_{e}$ is the concatenation of an injected instruction $s_{e}$ and injected data $x_{e}$, i.e., $p_{e}=s_{e}\oplus x_{e}$. Specifically, prompt injection attacks add a special string $z$ (called separator) between $x_{t}$ and $p_{e}$ to mislead the LLM into following the injected instruction instead of the target instruction. With an injected prompt, the LLM takes $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$ as input and generates a response that would accomplish an attacker-chosen injected task instead of the target task. Formally, $f(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})\approx f(s_{e}\oplus x_{e})$.
@@
According to a benchmark study (Liu et al., 2024a), Combined Attack is the most effective among these universal, heuristic-based attacks. Therefore, unless otherwise mentioned, we will use Combined Attack in our experiments to measure the vulnerability of LLM models under the attack of PoisonedAlign.
@@
Alignment:  LLMs are pre-trained on vast amounts of text data, and thus have the potential to generate harmful, biased, or misleading content if not properly aligned. Alignment aims to ensure that LLMs behave in ways that align with human values. Depending on the type of data used during alignment, there are two categories of alignment: 1) supervised fine-tuning (Wei et al., 2021) and 2) preference alignment (Ziegler et al., 2019; Ouyang et al., 2022; Christiano et al., 2017; Rafailov et al., 2023). In supervised fine-tuning, each alignment sample is a pair $(p,r)$, where $r$ is the desired response of an LLM for the given prompt $p$. In preference alignment, each alignment sample is a triple $(p,r_{1},r_{2})$, where the response $r_{1}$ is preferred over $r_{2}$, i.e., the LLM should be more likely to output $r_{1}$ than $r_{2}$ for the prompt $p$. Given the alignment data, supervised fine-tuning uses supervised learning to fine-tune the LLM, while preference alignment can be implemented using RLHF (Ziegler et al., 2019; Ouyang et al., 2022) or DPO (Rafailov et al., 2023).
@@
Poisoning alignment:  When the alignment data is collected from untrusted sources, e.g., third-party, crowd sourcing, or users, they may be poisoned by attackers (Baumgärtner et al., 2024; Pathmanathan et al., 2024; Wu et al., 2024; Wang et al., 2024b; Xu et al., 2024b; Rando and Tramèr, 2023). For instance, an attacker can poison the alignment data to embed a backdoor into the aligned LLM in supervised fine-tuning (Xu et al., 2024b). The backdoored LLM would generate specific, attacker-chosen responses when the prompt contains a backdoor trigger, such as a particular phrase. An attacker can also flip the preferences (i.e., $(p,r_{1},r_{2})$ is modified as $(p,r_{2},r_{1})$) in some alignment samples to reduce the performance of preference alignment (Pathmanathan et al., 2024). While PoisonedAlign is a data poisoning attack, it differs fundamentally from standard backdoor attacks. A conventional backdoor relies on a specific, discrete trigger, a keyword or phrase, to activate a specific, attacker-chosen behavior. In contrast, PoisonedAlign does not use a fixed trigger. Instead, it trains the model to become vulnerable to the general structure of a prompt injection. The ”trigger” is the pattern of a separator followed by new instructions, a pattern which can be instantiated in numerous ways. This makes the vulnerability more universal and potentially harder to detect and defend against than one tied to a static string.
@@
Our work is different from the above attacks in terms of both the attacker’s goals and the methods to create poisoned alignment samples. Previous studies have primarily considered prompt injection attacks that occur exclusively during inference. However, attackers can have strong incentives to carry out poisoning during the alignment stage, particularly when the attacker wishes to make inference-time prompt injection attacks more consistently effective. In particular, our attack goal is to poison alignment such that the aligned LLM is more vulnerable to inference-time prompt injection attacks while maintaining its foundational capability. Therefore, the primary novelty of PoisonedAlign lies not only in the mechanism of data poisoning itself, but also in its strategic goal: to degrade the model’s ability to maintain context and adhere to an initial instruction when a separator and a subsequent, conflicting instruction are introduced. It creates a widespread vulnerability rather than a narrow, trigger-based exploit. Due to the different attack goals, our attack requires a qualitatively different method to create poisoned alignment samples, compared to the above attacks that poison alignment.
3. Our PoisonedAlign
3.1. Threat Model
@@
  • Stealthiness:  The attack should be difficult to detect. This means the poisoned model $f_{p}$ must maintain its foundational capabilities, showing performance comparable to $f_{c}$ on standard evaluation benchmarks. While a single poisoned sample might appear suspicious in isolation, the attack is stealthy at scale, as detecting these patterns across massive datasets is a non-trivial data sanitation challenge.
@@
Another attack scenario involves LLM providers collecting the alignment data through crowd sourcing (Ryabinin and Gusev, 2020), where malicious crowd workers may provide poisoned alignment samples. Additionally, when an LLM (e.g., ChatGPT (OpenAI, 2024)) collects feedback/alignment data from its users, a malicious user (i.e., an attacker) can inject poisoned alignment data. Specifically, the attacker can query the LLM with a prompt $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$. When the LLM presents two response options collecting user feedback, the attacker selects the one that accomplishes its injected task. When the LLM only returns one response, the attacker can manipulate feedback–marking a response as “Bad” if it performs the target task, or as “Good” when it performs the injected task. The attacker can even ask the LLM to regenerate responses multiple times, repeating this feedback manipulation to potentially poison the alignment data further.
3.2. Creating Poisoned Alignment Samples
To create poisoned alignment samples, an attacker first collects a set of shadow prompt-response pairs $D_{s}=\{(p_{s_{i}},r_{s_{i}})\}_{i=1}^{N_{s}}$, where $r_{s_{i}}$ is a desired response for the prompt $p_{s_{i}}$. An attacker can collect $D_{s}$ either from question-answering benchmarks (Hendrycks et al., 2021; Rein et al., 2024; Cobbe et al., 2021) or through data synthesis (Xu et al., 2024a; Wang et al., 2024a). Then, our PoisonedAlign creates poisoned alignment samples based on $D_{s}$. In the following, we discuss creating poisoned samples for supervised fine-tuning and preference alignment separately.
@@ 3.2. Creating Poisoned Alignment Samples
4. Experiments
4.1. Experimental Setup
LLMs and alignment setting:  We use the following LLMs in our experiments: Llama-2-7b-chat (Touvron et al., 2023), Llama-3-8b-Instruct (Dubey et al., 2024), Gemma-7b-it (Mesnard et al., 2024), Falcon-7b-instruct (Almazrouei et al., 2023), and GPT-4o mini (OpenAI, 2024). The first four LLMs are open-source, while GPT-4o mini is closed-source. Unless otherwise mentioned, we apply DPO (Rafailov et al., 2023) to perform preference alignment. Since we cannot perform preference alignment for GPT-4o mini, we align it using Microsoft Azure’s supervised fine-tuning API in our ablation study.
@@ 4.1. Experimental Setup
Alignment datasets:  We use two popular alignment datasets in our experiments: HH-RLHF (Bai et al., 2022) and ORCA-DPO (Intel, 2024), which contain 169,352 and 12,859 alignment samples, respectively. Each alignment sample consists of a triple $(p,r_{1},r_{2})$. Due to computational constraints, we randomly sample 1,000 alignment samples from the training split of each dataset in our experiments. Note that supervised fine-tuning only uses the pairs $(p,r_{1})$ of each sample.
Poisoned alignment samples:  Given an alignment dataset $D$ (HH-RLHF or ORCA-DPO), for each alignment sample $(p,r_{1},r_{2})$, we obtain the pair $(p,r_{1})$; and these prompt-response pairs constitute our shadow dataset $D_{s}$. Then, we use PoisonedAlign to create poisoned alignment samples based on $D_{s}$. In our ablation study, we will also show that our attack is still effective when $D_{s}$ has no overlaps with $D$. Note that our poisoned alignment samples are constructed independently of the target or injected tasks used in the evaluation of prompt injection attacks. By default, we inject 10% of poisoned alignment samples into an alignment dataset, i.e., $|D^{\prime}|/|D|=10\%$, where $D^{\prime}$ is the set of poisoned alignment samples. We selected this rate because our experiments show it is near the point of saturation, allowing us to evaluate the maximum potential impact of the attack. The poisoned samples are constructed independently of the specific tasks used for the final evaluation.
Prompt injection attack:  Following Liu et al. (2024a), we adopt the following five prompt injection attacks in our experiments: Naive Attack (OWASP, 2023; Willison, 2022; Harang, 2023), Escape Characters (Willison, 2022), Context Ignoring (Perez and Ribeiro, 2022), Fake Completion (Willison, 2023), and Combined Attack (Liu et al., 2024a). Table 1 summarizes the separators used in these attacks. Unless otherwise mentioned, we use Combined Attack when creating poisoned alignment samples and in the evaluation of prompt injection attacks, since it demonstrated the highest effectiveness (Liu et al., 2024a). However, in our ablation study, we will also show that when the poisoned alignment samples are created using Combined Attack, the aligned LLM is also more vulnerable to other attacks.
Table 1. Prompt injection attacks and the corresponding separators. We use Combined Attack by default.
@@
Target and injected tasks:  Following Liu et al. (2024a), we use seven natural language tasks: duplicate sentence detection (DSD), grammar correction (GC), hate detection (HD), natural language inference (NLI), sentiment analysis (SA), spam detection (SD), and text summarization (Summ). We use MRPC dataset for duplicate sentence detection (Dolan and Brockett, 2005), Jfleg dataset for grammar correction (Napoles et al., 2017), HSOL dataset for hate content detection (Davidson et al., 2017), RTE dataset for natural language inference (Wang et al., 2019), SST2 dataset for sentiment analysis (Socher et al., 2013), SMS Spam dataset for spam detection (Almeida et al., 2011), and Gigaword dataset for text summarization (Rush et al., 2015). We use each task as either target or injected task. Therefore, we have $7\times 7$ target-injected task pairs. Unless otherwise mentioned, we use HD as the default target task and DSD as the default injected task. We use the target instruction $s_{t}$ and injected instruction $s_{e}$ for each task in Liu et al. (2024a). For completeness, we show them in Table 12 in the Appendix.
@@
Evaluation metrics:  To evaluate effectiveness of PoisonedAlign, we use Attack Success Value (ASV) (Liu et al., 2024a). Moreover, to evaluate stealthiness, we use accuracy to measure an LLM’s core capability on standard benchmarks. Specifically, we consider two variants of ASV: a loose version, ASVsoft, and a stricter version, ASVhard.  Liu et al. (2024a) used ASVsoft. Our evaluation metrics are defined as follows:
  • ASVsoft (Liu et al., 2024a):  ASVsoft defines a prompt injection attack as successful if the LLM completes the injected task correctly, regardless of whether it completes the target task. Given an LLM $f$, a dataset $D_{t}$ of target task samples $(p_{t},r_{t})$, a dataset $D_{e}$ of injected task samples $(p_{e},r_{e})$, and an attack that uses separator $z$, ASVsoft is formally defined as follows:
    (1) $\text{ASV}_{soft}=\sum\limits_{\begin{subarray}{c}(p_{t},r_{t})\in D_{t}\\ (p_{e},r_{e})\in D_{e}\end{subarray}}\frac{\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})}{|D_{t}||D_{e}|},$
    where $\mathcal{M}$ is the evaluation metric to measure whether the response $f(p_{t}\oplus z\oplus p_{e})$ matches the ground-truth answer $r_{e}$ of the injected task. For classification tasks like duplicate sentence detection, hate content detection, natural language inference, sentiment analysis, and spam detection, $\mathcal{M}$ is accuracy. In particular, $\mathcal{M}[a,b]$ is 1 if $a=b$ and 0 otherwise. For text summarization task, $\mathcal{M}$ is the Rouge-1 score (Lin, 2004), computed using the rouge 1.0.1 package with default settings. For grammar correction task, $\mathcal{M}$ is the GLEU score (Heilman et al., 2014), computed using the nltk 3.9.1 package with default settings.
@@
    (2) $\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})\cdot\mathbb{G}(f(p_{t}\oplus z\oplus p_{e})),$